OpenStudy (anonymous):

A pitched ball is hit by a batter at a 45degrees angle and just clears the outfield fence, 98m away. Assume that the fence is at the same height as the pitch and find the velocity of the ball when it left the bat. Neglect air resistance.

OpenStudy (jfraser):

use the equations of motion, in each direction, to solve backwards for Vx and Vy, then combine them using pythagorean theorem to find the original V

OpenStudy (jfraser):

\[\Delta s = V_{0}*t + \frac{1}{2}at^2\]\[V_f^2 = V_0^2 + 2a \Delta s\]\[V_f = V_0 + at\]

OpenStudy (anonymous):

You may use a relation that range (maximum horizontal displacement) of projectile is given by R = v^2 * sin(2A) / g Here A is the angle with the horizontal (here 45 ) and v is the velocity of projectile.(here the velocity with which the ball left the bat) .